OpenStudy (anonymous):

What is the excluded value for sqrt(4x+16)=x+1

OpenStudy (nikvist):

\[\sqrt{4x+16}=x+1\]\[4x+16\geq 0 \quad\wedge\quad x+1\geq 0\]\[x\geq -4 \quad\wedge\quad x\geq -1\quad\Rightarrow\quad x\geq -1\]

OpenStudy (nikvist):

excluded value is x<-1